OpenStudy (anonymous):

1/(x-1)(x-3),find out the domain and range

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

the domain is all the real numbers except x=1 and x=3 as you cannot divide by 0. the range is all the real numbers, as the function is unbounded. (for example near x=1 it takes big values)

OpenStudy (anonymous):

(x-1)(x-3) \[=x^2-4x+3\] Its range is (-D/4a,infinity) \[(-1,\infty)\] \[-1\le (x-1)(x-3) \le \infty\] \[\infty > \frac{1}{(x-1)(x-3)} \ge 0\] \[-1 \ge \frac{1}{(x-1)(x-3)} >-\infty\] Therefore the range is, \[R - (-1,0)\]

OpenStudy (anonymous):

Hey amogh... Is that a standard formula for finding range? (-D/4a,Infinity)

OpenStudy (anonymous):

For a quadratic equation, it is the minimum/maximum value (-b/2a,-D/4a) where D is the discriminant!